# Quasi-universal relations in the context of future neutron star detections

### v1

@@ II.3.2 Performing fits on agnostic equation of state sets
The usual approach for parameter extraction is to use fits. They correspond to a parametrized function (usually a polynomial) fitted to macroscopic parameters computed with a certain EoS set, and reported with a precision that attests to the largest difference between the fit and the set’s macroscopic parameters.
The pertinence of fits based on nuclear physics EoSs to represent the EoS variability in neutron stars must be questioned. First, the different core compositions have an impact on the modeling of macroscopic parameters. The first fits available were based on a limited number of EoS models. For example, the fits presented in Ref. [24] used three EoSs with purely nucleonic cores; the reported precision of $2\%$ for the relation $C(\Lambda)$ was shown in Ref. [33] to be too small when compared to several EoSs of various stiffness. Recognizing that the so called ”universal” relations were only quasi-independent of the description of neutron star’s interior, efforts were made to fit to larger sets of nuclear physics based EoSs, including various core compositions thus increasing the reported error of the fits; for example, in Ref. [5], the fits were performed using twenty EoSs with nucleonic cores, seven EoSs with hyperonic or kaon cores and three quark stars; a reported precision of 6.5% is given for neutron stars and 15% for quark stars. Neutron star’s core composition remains unknown, there is therefore no reason to perform the fit on sets containing more nucleonic models than hybrid or hyperonic ones. This is however what is done in practice, systematically with fits based on nuclear physics EoS, likely because nucleonic cores are easily accessible in the literature. A similar reasoning applies to the nuclear physics approach used to establish the set of EoSs on which the fits are based on: relativistic mean field based models tend to be stiffer than e.g., Skyrme energy density functional models. Secondly, there is no reason to reduce the set of EoSs to known hypothesis of the core composition nor our approaches of nuclear physics. To that end, the more relevant set of EoSs to perform the fits should be agnostic, as presented e.g. in Ref. [25, 23].
In this paper, we perform fits of the three relations presented in Section II.1.1 with the different EoS sets presented in this paper. For each relation and each set, a non-linear least square method is used to determine the parameter $a_{k}$ with $k\in[0,5]$ in
@@
An error is associated to each fit and denoted ${(\Delta X_{\rm fit})^{\rm S}}$ with $X$ the fitted quantity and $S$ the set of EoSs. The fits are performed using macroscopic parameters of neutron star with at least a mass of 1.0 ${}_{\odot}$ up to the last stable mass configuration. The error is assessed from the full range of the (finite) nuclear set, and the 99% percentiles of the GP+astro set and MM+$\chi$+PSR set. The fit parametrizations for $C(\Lambda)$, $\bar{I}(\Lambda)$ and $\bar{I}(C)$ presented in Ref. [5] (later denoted Yagi & Yunes) are also used as a comparison.
@@ II.3.3 Marginalization of the error
To include the fit error ${(\Delta X_{\rm fit})^{\rm S}}$ in the extraction of a parameter, we must specify a marginalization technique. This can be done by defining an extra variable $\delta X$ with a random distribution reflecting the fit uncertainty, such that each sample fitted parameter is multiplied by a factor $(1+\delta X_{\rm fit}^{S})$. In this paper, we describe this distribution in relation to the error on the fit for a given relation ${(\Delta X_{\rm fit})^{\rm S}}$. For example, in Refs. [11, 47], the marginalization consisted of drawing the fit error parameter $\delta X$ from a Gaussian function peaked on the fitted relation with a ${(\Delta C_{\rm fit})^{\rm yy}}$ variance; with the fit’s error assigned to a 3-$\sigma$ precision, the standard deviation for the Gaussian distribution is ${(\Delta X_{\rm fit})^{\rm S}/3}$. We refer to this technique as the Gaussian marginalization.
In the absence of information on the distribution of macroscopic parameters computed from the EoS set used to perform the fit, we propose that users choose a uniform distribution instead of a Gaussian one: ${\delta X_{\rm fit}^{S}\in[-(\Delta X_{\rm fit})^{\rm S};(\Delta X_{\rm fit})^{\rm S}]}$ is randomly drawn from a uniform distribution; this approach is referred to later on as the uniform marginalization. It leads to a larger systematic error on the extraction of the macroscopic parameter but avoids the fit line bias. To illustrate that the distribution of points is not necessarily a Gaussian for a given set of EoS, we present the distribution of compactness and dimensionless moment of inertia around the fit for equation of state sets in Appendix B.
One can also assess the full extent of the fit error. We take the two extremes of the relation ${X_{\rm fit\pm}^{S}(Z)=X_{\rm fit}^{S}(Z)\pm(\Delta X_{\rm fit})^{\rm S}}$. We then perform the parameter extraction and obtain two distributions which represent the two extremes of the fit error; we refer to this approach as the fit limits.
II.4 Equation of state inference
@@
Astrophysical sources inform the EoS: sources such as the simulated ones presented Section II.3.1 therefore have the potential to constrain the sets of EoSs presented in Section II.2. We use the following EoS inference approaches.
For the X-ray simulation, we use a simple likelihood estimation on the 1-$\sigma$ STROBE-X sensitivity data and the MM+$\chi$+PSR and GP+astro sets. For the gravitational wave simulated data, we use the EoS inference technique discussed in Refs. [48, 40, 49, 50] with the software Likelihood Weighting Protocol (LWP [^5] https://git.ligo.org/reed.essick/lwp) to infer on the MM+$\chi$+PSR and GP+astro set with the two neutron stars of the binary coalescence in Cosmic Explorer and Einstein Telescope sensitivity.
With this newly informed sets of EoSs, we can extract parameters from the EoS $M(R)$ and $M(I)$ and compare the EoS dependent parameter extraction to data-driven relation technique.
@@
Figure 2: $M(R)$ extracted from the $C(\Lambda)$ relation and a simulated GW signal in the sensitivity of O4 (orange) and Cosmic Explorer and Einstein telescope (green); contours are shown at 1-$\sigma$. Results are presented for the fit line (plain lines), with the Gaussian marginalization of the error (dashed line), and with the uniform marginalization of the error (dotted line); the limits of the error are presented in shades of colors.
@@ III.1 Comparison of the different equation of state sets
In Fig. 1 we present the relations of Section II.1.1 established from different sets of EoS: the nuclear set, the GP+astro and the MM+$\chi$+PSR set. In black is also presented the fits of Yagi & Yunes.
@@
The agnostic GP+astro set presents the largest EoS variability. It explores a larger astrophysics parameter space than the nuclear set and the agnostic MM+$\chi$+PSR set, and overlaps them both, thus showing that the chiral effective field theory constraints used for the meta-modeling approach offer stronger constraints than current astrophysical measurements from NICER and GW170817. We note also that the GP+astro set does not encompass the meta-model (nor the nuclear) set in a symmetric way, particularly at low compactness: the meta model set favors higher tidal deformabilities and moment of inertia, in other words, stiffer EoSs.
@@
The nuclear set and the meta model do not perfectly overlap, as some of the EoSs used in the nuclear set are not in accordance with chiral effective field theory constraints. The nuclear set is in good accordance with the fit of Yagi & Yunes, which was also based on complete nuclear physics calculation EoSs. The nuclear set was established using 61 EoSs with various core compositions (including deconfined quarks), however it presents a stronger EoS invariance than the Yagi & Yunes fit based on $\sim 30$ EoSs. For example, the error associated to the Yagi & Yunes fit of $C(\Lambda)$ is of 6.5% and 15% excluding and including quark stars respectively. Our fit for this relation (see Appendix for parameters) gives a 6% maximum error (see Table 1 in the Appendix) even though it includes a significant portion of hybrid models. The selection of modern EoS which are calibrated to astrophysical and nuclear physics data plays a role in the smaller error (although not significant) of our fit: for example, the set used to establish the Yagi & Yunes fit includes models presenting low radii at fixed mass (i.e. high tidal deformability and moment of inertia), which microphysics parameters have been reproved by nuclear physics laboratory experiments e.g. WFF1 and WFF2 [51].
@@
The relation $C(\Lambda)$ in Fig. 0(a) and $\bar{I}(C)$ in Fig. 0(b) present a clear EoS variability. The relation $\bar{I}(\Lambda)$ in Fig. 0(c) is EoS invariant to the point that we cannot distinguish differences between the sets of EoSs: this relation is sufficiently strongly universal to overcome the EoS variability emerging from the use of different EoS sets. This result is in accordance with Ref. [23] which quantifies the degree of universality of various relations: we also find sub-percent errors on the fits (see Table 1 in the Appendix). Contrary to the compactness involved relations in which the different sets and their various EoS constraints can be distinguished, the $\bar{I}(\Lambda)$ universality cannot be attributed to our use of data driven EoS sets.
The particular EoS invariance of the relation $\bar{I}(\Lambda)$ was discussed in Ref. [5] and reference therein in the larger context of I-Love-Q relations. One of the hypothesis put forward to explain the EoS invariance of I-Love-Q relations was the lack of variability far from the core (low densities). The GP+astro set presents a unique crust while the meta model offers crust variability yet the former presents a slightly larger error on the fit of $\bar{I}(\Lambda)$ than the latter: we did not observe an obvious loss of invariance when including the nuclear physics constrained crust variability of the meta model.
@@ III.2.1 Radius from binary neutron star mergers
In Fig. 2, we present the extraction of the radius from the simulated gravitational wave signal of a binary neutron star merger using the relation $C(\Lambda)$; the O4 sensitivity is presented in orange and the Cosmic Explorer and Einstein telescope sensitivity is presented in green. Fig. 1(a) used the fit based on the meta-modeling set of EoS while Fig. 1(b) used the Gaussian process set.
The radius extraction using the Gaussian or the uniform marginalization of the error, or simply the fit line and without including any error, are (roughly) equivalent at O4 sensitivity. This fit limits (in shaded colors) are larger at that sensitivity but not significantly relative to the size of the contour, although it is more pronounced for the Gaussian process fit.
In the case of 3rd generation sensitivity, we first see that there is a significant difference between the Gaussian and uniform marginalization, as well as with the fit line without inclusion of the error: the Gaussian marginalization has overestimated the precision on the radius by around a factor of two compared to uniform marginalization. We can conclude that the technique used to include the error has an impact on the results in this sensitivity of the detectors. We also see on that figure that the extraction of the radius from the fit line is significantly thinner than the error marginalization: this indicates that the parameter extraction contours at that sensitivity are related to the fit error and not the the tidal deformability recovery from the gravitational wave signal. This is confirmed by the limits of the error presented in shades of green: they are at the opposite ends of the uniform marginalization and do not overlap at all.
Finally, the meta-model (MM+$\chi$+PSR) fit offer a better precision on the parameter extraction than the Gaussian process (GP+astro) fit, as the maximum error presented in Table 1 indicated. This indicates that in parameter extraction, current astrophysical constraints are not equivalent to nuclear theory calculations in data-driven relations. We also note that higher radii are extracted for a given mass with the Gaussian process fit, in accordance with our discussion of Fig. 0(a) in Section III.1.
Figure 3: $I(M)$ extracted from the $\bar{I}(\Lambda)$ relation and a simulated GW signal in the sensitivity of O4 (orange) and Cosmic Explorer and Einstein telescope (green); contours are shown at 1-$\sigma$. Results are presented for the fit line (plain lines), with the Gaussian marginalization of the error (dashed line), and with the uniform marginalization of the error (dotted line); the limits of the error are presented in shades of colors. Fits used are that of GP+astro.
@@ III.2.2 Moment of inertia from binary neutron star mergers
Figure 4: $I(M)$ extracted from the $\bar{I}(C)$ relation and the simulated X-ray signal in the sensitivity of NICER (orange) and Strobe-X (green); contours are shown at 1-$\sigma$. Results are presented for the fit line (plain lines), with the Gaussian marginalization of the error (dashed line), and with the uniform marginalization of the error (dotted line); the limits of the error are presented in shades of colors.
In Fig. 3, we present the extraction of the moment of inertia using the relation $\bar{I}(\Lambda)$ and the fit based on the Gaussian process set (which produces the largest fit error, see Table 1). In that case, we can see that the two marginalization techniques, the limit of the error and the extraction from the fit line all overlap, both for O4 sensitivity and CE+ET sensitivity. This is not explained by the parameter estimation of the tidal deformability, as we use the same source and same set of EoS as for Fig. 2 but by the nature of the quasi-universality of the relation $\bar{I}(\Lambda)$. As can be seen in Fig. 0(c), this relation presents a significant universality, which makes considerations of the marginalization technique, or even purely of the error, irrelevant. The error associated to the parameter estimation dominates whatever the sensitivity of the detector. We conclude that the universality of the relation $\bar{I}(\Lambda)$ will strongly hold with the 3rd generation of gravitational wave detectors.
@@ III.2.3 Moment of inertia from X-ray detections
In Fig. 4, we present the extraction of the moment of inertia using the relation $\bar{I}(C)$. For the meta-model fits presented in Fig 3(a), the marginalization technique does not impact the results. For the fits based on the Gaussian process set presented in Fig. 3(b), the results are impacted at high sensitivity (STROBE-X) as the Gaussian marginalization underestimates the error by almost a third compared to the uniform marginalization. The constraints brought forth by nuclear physics in the meta-modeling allow us to consider that STROBE-X sensitivity still dominates (or is at least equivalent) to the EoS variability of the meta-model based data-driven relation $\bar{I}(C)$.
III.3 Equation of state inference vs data-driven relations

### v2

@@ II.3.2 Performing fits on agnostic equation of state sets
The usual approach for parameter extraction with quasiuniversal relations is to use parametrized functions (usually a polynomial) fitted to macroscopic parameters computed with a set of EoS, and reported with a precision that attests to the largest difference between the fit and the set’s macroscopic parameters.
The first fits available were based on a limited number of EoS models. For example, the fits presented in Ref. [36] used three EoSs with purely nucleonic cores; the reported precision of $2\%$ for the relation $C(\Lambda)$ was shown in Ref. [46] to be too small when compared to several EoSs of various stiffness. Recognizing that the so called ”universal” relations were only quasi-independent of the description of neutron star’s interior, efforts were made to fit to larger sets of nuclear physics based EoSs, including various core compositions thus increasing the reported error of the fits; for example, in Ref. [10], the fits were performed using 20 EoSs with nucleonic cores, seven EoSs with hyperonic or kaon cores and three quark stars; a reported precision of 6.5% is given for neutron stars and 15% for quark stars. As a neutron star’s core composition remains unknown, a bias may be introduced by performing the fit on sets containing more nucleonic models than hybrid or hyperonic ones. Bias may also be introduced by the nuclear physics approach used for the set of EoSs on which the fits are based: relativistic mean field based models tend to be stiffer than, e.g., Skyrme energy density functional models. To mitigate the impact of these choices, we choose sets of EoSs that are more agnostic, as has been done, e.g., in Ref. [37, 35, 58].
In this paper, we fit each of the three relations presented in Sec. II.1.1 with the different EoS sets considered in this paper. For each relation and each set, a nonlinear least square method is used to determine the parameter $a_{k}$ with $k\in[0,5]$ in
@@
The fits are performed using macroscopic parameters of neutron stars with at least a mass of 1.0 M up to the last stable mass configuration. An error range is associated to each fit and denoted ${(\Delta X_{\rm fit})^{\rm S}}$ with $X$ the fitted quantity and $S$ the set of EoSs. The error is assessed from the full range of the (finite) nuclear set, and the 99% percentiles of the GP+astro set and MM+$\chi$+PSR set. For the GP+astro and MM+$\chi$+PSR sets, there are no defined edge to the range because the EoS are drawn from an underlying distribution that can, in principle, extend to infinity. The fit parametrizations for $C(\Lambda)$, $\bar{I}(\Lambda)$, and $\bar{I}(C)$ presented in Ref. [10] (later denoted Yagi & Yunes) are also used as a comparison.
@@ II.3.3 Marginalization of the error
To include the fit error ${(\Delta X_{\rm fit})^{\rm S}}$ in the extraction of a parameter, we specify a marginalization over this systematic uncertainty, as seen previously in e.g. Refs. [33, 59, 40]. This is done by introducing an extra parameter $\delta X$ with a random distribution reflecting the fit uncertainty. Each extracted parameter is multiplied by a factor $(1+\delta X_{\rm fit}^{S})$. For example, when applying $C_{\text{fit}}$ as in Refs. [33, 59], the marginalization consists of drawing the fit error parameter $\delta X$ from a Gaussian function peaked at zero with standard deviation ${(\Delta C_{\rm fit})^{\rm S}/3C_{\text{fit}}}$, so that results are within ${\Delta C_{\rm fit}}$ at 3-$\sigma$. We refer to this technique as the Gaussian marginalization.
In the absence of information on the distribution of macroscopic parameters computed from the EoS set used to perform the fit, we propose that users choose a uniform distribution instead of a Gaussian one: ${\delta X_{\rm fit}^{S}\in[-(\Delta X_{\rm fit})^{\rm S};(\Delta X_{\rm fit})^{\rm S}]}$ is randomly drawn from a uniform distribution; we refer to this approach as uniform marginalization. It leads to a larger error on the extraction of the macroscopic parameter but avoids preferring the exact fit. To illustrate that the distribution of points is not necessarily a Gaussian for a given set of EoS, we present the distribution of compactness and dimensionless moment of inertia around the fit for equation of state sets in Appendix C.
We also show the extremes of the quasiuniversal relation as a way to represent the systematic uncertainty. We consider ${X_{\rm fit\pm}^{S}(Z)=X_{\rm fit}^{S}(Z)\pm(\Delta X_{\rm fit})^{\rm S}}$. We perform the parameter extraction at each extreme of the range and obtain two distributions which represent the two extremes of the fit error. We refer to this approach as the fit limits.
@@
Figure 2: $M(R)$ extracted from the fit of $C(\Lambda)$ and a simulated GW signal in the sensitivity of O4 (orange) and CE and ET (green); contours are shown at 1-$\sigma$. Results are presented for the exact fit (plain lines), with the Gaussian marginalization of the error (dashed line), and with the uniform marginalization of the error (dotted line); the limits of the error are presented in shades of colors.
@@ III.1 Comparison of the different equation of state sets
In Fig. 1 we present the relations of Sec. II.1.1 established from different sets of EoS: the nuclear set, the GP+astro and the MM+$\chi$+PSR set. In black is also presented the fits of Yagi & Yunes.
@@
The agnostic GP+astro set presents the largest EoS variability. It explores a larger space than the nuclear set and the agnostic MM+$\chi$+PSR set, and overlaps them both, showing that the chiral effective field theory constraints used in the metamodeling approach offer stronger constraints than current astrophysical measurements from NICER and GW170817. We note also that the GP+astro set does not encompass the metamodel (nor the nuclear) set in a symmetric way, particularly at low compactness: the meta model set favors higher tidal deformabilities and moment of inertia, in other words, stiffer EoSs.
@@
The nuclear set and the meta model do not perfectly overlap, as some of the EoSs used in the nuclear set are not in accordance with chiral effective field theory constraints. The nuclear set is in good accordance with the fit of Yagi & Yunes, which was also based on complete nuclear physics calculation EoSs. The nuclear set was established using 61 EoSs with various core compositions (including deconfined quarks), however it has stronger EoS invariance than the Yagi & Yunes fit based on $\sim 30$ EoSs. The error associated to the Yagi & Yunes fit of $C(\Lambda)$ is of 6.5% and 15% excluding and including quark stars, respectively. Our fit for this relation gives a 6% maximum error (see Table 1 in the Appendix D) even though it includes a significant portion of hybrid models. The smaller error is related to the selection of modern EoSs which are calibrated to astrophysical and nuclear physics data: for example, the set used to establish the Yagi & Yunes fit includes models with low radii at fixed mass (i.e., high tidal deformability and moment of inertia), for which microphysics parameters have been disfavored by nuclear physics laboratory experiments, e.g., WFF1 and WFF2 [60].
@@
The relation $C(\Lambda)$ in Fig. 1(a) and $\bar{I}(C)$ in Fig. 1(b) retain a clear EoS variability. The relation $\bar{I}(\Lambda)$ in Fig. 1(c) is EoS invariant to the point that we cannot distinguish differences between the sets of EoSs: this relation is sufficiently universal to overcome the EoS variability emerging from the use of different EoS sets. This result is in accordance with Ref. [35] which quantifies the degree of universality of various relations: we also find subpercent errors on the fits (see Table 1 in the Appendix D).
@@ III.2 Error marginalization
In this section, we compare different error marginalization techniques for quasiuniversal relation applications as described in Sec. II.3.3 in the context of current and future detections. For gravitational wave data (see Sec. II.3.1), simulated $M$ and $\Lambda$ distributions are used in the fits of $C(\Lambda)$ and $\bar{I}(\Lambda)$, with different error marginalization, to obtain distributions of $R$ and $I$ respectively. For x-ray data (see Sec.. II.3.1), simulated $M$ and $R$ distributions are used in the fit of $\bar{I}(C)$, with different error marginalization, to obtain distributions of $I$.
@@ III.2.1 Radius from binary neutron star mergers
In Fig. 2, we present the extraction of the radius from the simulated gravitational wave signal of a binary neutron star merger using the fit of the quasiuniversal relation $C(\Lambda)$; the O4 sensitivity is presented in orange and the CE and ET sensitivity is presented in green. Fig. 2(a) used the fit based on the metamodeling set of EoS while Fig. 2(b) used the Gaussian process set.
The radius extraction using the Gaussian or the uniform marginalization of the error, or simply the fit line and without including any error, are (roughly) equivalent at O4 sensitivity. These fit limits (in shaded colors) are larger at that sensitivity but comparable to or smaller than to the size of the contour.
In the case of third generation sensitivity, we first see that there is a significant difference with the fit line without inclusion of the error as well as between the Gaussian and uniform marginalization: the Gaussian marginalization has overestimated the precision on the radius by around a factor of 2 compared to uniform marginalization. We can conclude that the technique used to include the systematic error has an impact on the results at this sensitivity. We also see on that figure that the range of the radius extracted with the fit line is significantly smaller than after error marginalization: this indicates that the parameter extraction contours at that sensitivity are related to the fit error and not the the tidal deformability recovery from the gravitational wave signal. This is confirmed by the fit limits presented in shades of green: they are at the opposite ends of the uniform marginalization and do not overlap at all.
Finally, the metamodel (MM+$\chi$+PSR) fit offers a better precision on the parameter extraction than the Gaussian process (GP+astro) fit, as the maximum error presented in Table 1 indicated. This indicates that in parameter extraction, current astrophysical constraints are less constraining than nuclear theory calculations. We also note that higher radii are extracted for a given mass with the Gaussian process set, in accordance with our discussion of Fig. 1(a) in Sec. III.1.
Figure 3: $I(M)$ extracted from the fit of $\bar{I}(\Lambda)$ and a simulated GW signal in the sensitivity of O4 (orange) and CE and ET (green); contours are shown at 1-$\sigma$. Results are presented for the fit line (plain lines), with the Gaussian marginalization of the error (dashed line), and with the uniform marginalization of the error (dotted line); the fit limits are presented in lighter shades. Fits used are that of GP+astro.
@@ III.2.2 Moment of inertia from binary neutron star mergers
Figure 4: $I(M)$ extracted from the $\bar{I}(C)$ relation and the simulated x-ray signal in the sensitivity of NICER (orange) and STROBE-X (green); contours are shown at 1-$\sigma$. Results are presented for the fit line (plain lines), with the Gaussian marginalization of the error (dashed line) and with the uniform marginalization of the error (dotted line); the fit limits are presented in lighter shades.
In Fig. 3, we present the extraction of the moment of inertia using the fit of $\bar{I}(\Lambda)$ based on the Gaussian process set (which produces the largest fit error, see Table 1). In that case, we can see that the two marginalization techniques, the limit of the error and the extraction from the fit line all overlap, both for O4 sensitivity and CE+ET sensitivity. We use the same source and same set of EoS as for Fig. 2, so this overlap results from the quasiuniversality of the relation $\bar{I}(\Lambda)$. As can be seen in Fig. 1(c), this relation has stronger universality, which makes considerations of the marginalization technique, or even purely of the error, irrelevant. The statistical uncertainty associated with the parameter estimation dominates whatever the sensitivity of the detector. We conclude that the universality of the relation $\bar{I}(\Lambda)$ will hold with the third generation of gravitational wave detectors.
@@ III.2.3 Moment of inertia from x-ray detections
In Fig. 4, we present the extraction of the moment of inertia using the relation $\bar{I}(C)$. For the metamodel fits presented in Fig 4(a), the marginalization technique does not impact the results. For the fits based on the Gaussian process set presented in Fig. 4(b), the results are impacted at high sensitivity (STROBE-X) as the Gaussian marginalization underestimates the error by almost a third compared to the uniform marginalization. The constraints brought forth by nuclear physics in the metamodel allow us to consider that STROBE-X sensitivity still dominates (or is at least equivalent) to the EoS variability of the metamodel based quasiuniversal relation $\bar{I}(C)$.
III.3 Quasiuniversal relations vs EoS inference
In this section, we compare the extraction of the radius and the moment of inertia using two methods: the quasiuniversal fits and EoS inference. We consider a simulated CE+ET gravitational wave detection recovering $M$ and $\Lambda$ and a simulated STROBE-X detection recovering $M$ and $R$, following the description of Sec. II.3.1. For quasiuniversal relations, we use the distributions of $R$ and $I$ presented in Fig. 2, Fig. 3, and Fig. 4 in the case of uniform marginalization.
For EoS inference, we use the simulated data to infer constraints on the EoS with a simple likelihood estimation on the 1-$\sigma$ data and the MM+$\chi$+PSR and GP+astro sets. We then use the updated EoS distribution and general relativity to compute $R$ and $I$ from the mass distributions of the of the sources using the methods of Sec. II.